OpenStudy (sleepyjess):

An auditorium has 30 rows with 10 seats in the first row, 12 in the second row, 14 in the third row, and so forth. How many seats are in the auditorium?

Nnesha (nnesha):

\[a_1 = 10 ~~~~a_2 =12\] do you know the common difference ?

Nnesha (nnesha):

it's Arithmetic seq bec first one is 10 then 2nd one is 12 3rd one is 14

Nnesha (nnesha):

do you know the formula ??

OpenStudy (aaronandyson):

Formula for the sum of arithmetic sequence is Sn = (n/2) (A1 + An) where A1 is the first and An is the last term. In our case, sequence has 20 terms (20 rows in auditorium) and the last term (= number of seats in 20th row) is A20 = 2(20 + 4) = 48 now the sum is S20 = (20/2) (10 + 48) = 10 * 58 = 580